How many gallons of 50% solution must be mixed with 100 Gallons of 20% solution of salt that is 40% salt
Answer by jorel1380(3719)

m(.5)+100(.2)=(m+100).4
.5m+20=.4m+40
.1m=20
m=20/.1=200
You need 200 gallons of 50% solution for your mix..